# A Chinese restaurant process for multiset permutations

### v2

@@ 1 Introduction
A growth rule which received considerable attention in the literature is the following preferential attachment algorithm on permutations written as the product of cycles known as the Chinese restaurant process (CRP) [1, 6, 9, 14, 17]. For $n\geq 1$, given a permutation of $n-1]:=\left\{1,2,\ldots,n-1\right\}$ which has been constructed at step n1n-1, element $n$ is either appended to the permutation as a new singleton cycle with probability $\frac{\theta}{\theta+n-1}$, or inserted into a random position within the existing permutation with probability equal to $\frac{n-1}{\theta+n-1}$. Let $S_{n}$ be the set of permutations on $[n]$. The permutation obtained at step $n$ has the Ewens distribution, which is the uniform distribution on $n!$ permutations in the case $\theta=1$, and, in general, has probability distribution $\theta^{|\pi|}/\theta_{(n)}$, $\pi\in S_{n}$, where $\theta_{(n)}:=\theta(\theta+1)\cdots(\theta+n-1)$ and $|\pi|$ is the number of cycles in $\pi$.
Let $\Pi_{n}$ be the random permutation of $[n]$ at the $n$th step of the CRP. The distribution of $\Pi_{n}$ is invariant under conjugations. For different orders the permutations are consistent, in the sense that $\Pi_{m}$, for $m<n$, can be derived from $\Pi_{n}$ by removing elements $m+1,\ldots,n$ from their cycles and deleting empty cycles if necessary.
@@
$\pi=(x_{11}\ldots x_{1n_{1}}y_{1})(x_{21}\ldots x_{2n_{2}}y_{2})\cdots(x_{t1}\ldots x_{tn_{t}}y_{t}),\quad t\geq 0,$
@@
which corresponds to the distribution of the sum of $n$ independent Bernoulli variables $K_{n}=\sum_{i=1}^{n}I_{i}$, $I_{i}\sim{\rm Bernoulli}(i/(i+\theta-1))$. We have $I_{i}=1$ exactly when the element $i$ starts a new cycle. As $n\to\infty$,
$K_{n}\sim\theta\log n~~{\rm a.s.},~~~\frac{K_{n}-\theta\log n}{\sqrt{\theta\log n}}\stackrel{{\scriptstyle d}}{{\to}}{\rm N}(0,1).$
In this paper, a CRP will be defined for multiset permutations will be derived using the methods in [17]. Multiset permutations are of fundamental interest in combinatorics [16] and random multiset permutations have been studied [7].
@@ 2 Multiset permutations and their factorisations into
There are at least two different ways of defining cycles of multiset permutations [13, 16]. Knuth’s way [13] will be used in this paper. In this section, we summarise the discussion and results on multiset permutation factorisation taken from Section 5.1.2 of [13] that is needed in this paper. In particular, we are going to use Theorem 1 below to factor multiset permutations. A multiset is like a set except that it can have repetitions of identical elements, for example
@@
The intercalation product $\pi_{1}\perp\pi_{2}$ of two multiset permutations $\pi_{1}$ and $\pi_{2}$ is defined in [4, 8] and is obtained by
@@
For example
3 1 4 1 2 2 4 4 1 4=3 1 2 4 4 1 2 4 1 43\ 1\ 4\ 1\ 2\ $\perp\$2\ 4\ 4\ 1\ 4=3\ 1\ 2\ 4\ 4\ 1\ 2\ 4\ 1\ 4
since
$\left(\begin{array}[]{c c c c c c}1&1&2&3&4\\ 3&1&4&1&2&\end{array}\right)\perp\left(\begin{array}[]{c c c c c}1&2&4&4&4\\ 2&4&4&1&4\end{array}\right)$
@@
and so (1) is a cycle. For these general cycles, $(x_{1}x_{2}\ldots x_{n})$ is not always the same as $(x_{2}\ldots x_{n}x_{1})$.
Knuth [13] proves the following factorisation of multiset permutations into cycles.
Theorem 1 ([13]).
@@
$\pi=(x_{1,1}\ldots x_{1,n_{1}}y_{1})\perp(x_{2,1}\ldots x_{2,n_{2}}y_{2})\perp\cdots\perp(x_{t-1,1}\ldots x_{t-1,n_{t-1}}y_{t-1})$ (6)
for $t\geq 1$, where the following two conditions are satisfied:
@@
    $y_{i}<x_{i,j}{\rm\ for\ }1\leq j\leq n_{i},\,1\leq i\leq t-1$.
For example, the factorisation of (1) corresponding to Theorem 1 is
[…]
@@ 3 The CRP for multiset permutations
where $N_{t}=n_{1}+\cdots+n_{t}$. Note that $S_{n}=S(\underbrace{{1,1,\ldots,1}}_{\text{$n$ times}}).$ The random multiset permutation $\tilde{\Pi}_{t}$ of our process at the $t$th step is an element of $S(n_{1},\ldots,n_{t})$, represented in accordance with (6). At step $t\in[s]$, there are $n_{t}$ copies of $t$ either to be inserted in the existing permutation $\tilde{\Pi}_{t-1}$ or to be put into singleton cycles. (It is not possible to put more than one copy of $t$ in a new cycle because of condition 2 of Theorem 1.) Suppose that $k_{t}\in\{0\}\cup[n_{t}]$ of the $n_{t}$ copies of $t$ start $k_{t}$ new singleton cycles containing $t$, and the $n_{t}-k_{t}$ other copies of $t$ are inserted into the existing permutation. The $n_{t}-k_{t}$ copies of $t$ can be inserted to the left of any of the $y_{i}$ or $x_{ij}$ in (6). A ‘stars and bars’ argument shows that the number of ways of inserting $n_{t}-k_{t}$ copies of $t$ to the left of an element in the multiset $S(n_{1},\ldots,n_{t-1})$ formatted like (6) is
@@
$F_{t}(1)=\sum_{k_{t}=0}^{n_{t}}\binom{N_{t-1}-1+n_{t}-k_{t}}{n_{t}-k_{t}}=\binom{N_{t}}{n_{t}}$, (8)
results from a well known identity; see page 161 of [10]. Note that (8) is the total number of ways of extending the permutation from time t1t-1 to time $t$. At time $t\in[s]$, $k_{t}$ new singleton cycles containing $t$ are created and the other $n_{t}-k_{t}$ copies are inserted into the existing permutation in a specified way with probability $\theta^{k_{t}}/F_{t}(\theta)$. If $n_{t}=1$ for all $t$, then $F_{t}(\theta)=\theta+t-1$ and we recover the usual CRP. An induction argument on $t$ shows that all possible $\pi\in S(n_{1},\ldots,n_{t})$ can be constructed in this manner. This can also be seen from the identity
@@
We will use the method in [17]. An arborescence is a directed, rooted tree with edges oriented in agreement with paths from the root to the leaves. Our arborescence has vertex set $V=\bigcup_{s=0}^{t}\Pi(n_{1},\ldots,n_{s})$ and the root is $\emptyset$ (when $s=0$). The leaves of the arborescence are $\Pi(n_{1},\ldots,n_{t})$. Give each leaf $\pi\in\Pi(n_{1},\ldots,n_{t})$ the weight $\theta^{|\pi|}$, where $|\pi|$ is the number of cycles in $\pi$. There is a directed edge from $\pi_{1}\in\Pi(n_{1},\ldots,n_{s-1})$ to $\pi_{2}\in\Pi(n_{1},\ldots,n_{s})$, $s\in[t]$, if and only if, for some $k_{s}\in\{0\}\cup[n_{s}]$, $\pi_{2}$ is obtained from $\pi_{1}$ by inserting $n_{s}-k_{s}$ copies of $s$ to the left of elements of $\pi_{1}$, when $\pi_{1}$ is written as (1), and creating $k_{s}$ singleton cycles each containing $s$. For each $\pi\in V$, define $Q(\pi)$ to be the sum of the weights of all leaves $\pi^{\prime}\in\Pi(n_{1},\ldots,n_{t})$ for which there is a directed path from $\pi$ to $\pi^{\prime}$. For $\pi\in\Pi(n_{1},\ldots,n_{s})$, we have
@@
These are the transition probabilities of the CRP. Theorem 1 of [17] states that if we put transition probabilities (11) on the edges of the arborescence, and let the leaves be absorbing states, then we obtain a Markov chain whose probability of absorption on a given leaf is proportional to the weight of the leaf. By our choice of weight we are done. ∎
@@ 4 The distribution of the number of cycles
Let $X_{t}$ denote the number of new cycles added in moving from time t1t-1 to time $t$, $t\geq 1$, which is the choice of $k_{t}$ above. The $X_{t}$ are independent and have p.g.f.
@@
It is easily checked that (13) is (14) with $N=N_{t}$< $M=N_{t}-n_{t}=N_{t-1}$, $r=1$ and $\kappa=k+1$. The expectation of $Y$ is
@@
and the third centred moment (see [11, 12]) is
@@
It follows that the number of cycles $K_{t}$ has expectation
${\mathbb{E}}(K_{t})=\sum_{s=1}^{s}\frac{n_{s}}{N_{s-1}+1}$
@@
Note that $n_{t}$ only appears in the numerators in (15) and (16), and therefore, by increasing $n_{t}$, with $n_{1},n_{2},\ldots,n_{t-1}$ fixed, we can make the expectaion and variance as large as desired.
Given sequences $a_{t}$ and $b_{t}$, we use the notation $a_{t}=O(b_{t})$ to mean $a_{t}\leq Cb_{t}$ for a constant $C>0$; $a_{t}\asymp b_{t}$ to mean $C_{1}b_{t}\leq a_{t}\leq C_{2}b_{t}$ for constants $C_{1}>0$, $C_{2}>0$; and $a_{t}\sim b_{t}$ to mean $\lim_{t\to\infty}a_{t}/b_{t}=1$. We will put the bound the $n_{t}=O(1)$ to give a central limit theorem for $K_{s}$ for all $\theta$.
@@
${\mathbb{E}}(K_{t})\asymp\log t\quad{\rm Var}(K_{t})\asymp\log t,$ (20)
@@
$K_{n}\sim{\mathbb{E}}(K_{n})~~{\rm a.s.}~~~~~~\frac{K_{t}-{\mathbb{E}}(K_{t})}{\sqrt{\rm Var}(K_{t})}\stackrel{{\scriptstyle d}}{{\to}}{\rm N}(0,1).$ (21)
@@
First we prove the theorem for $\theta=1$. From (15), (16) and (19) we get ${\mathbb{E}}(X_{t})\asymp 1/t$ and ${\rm Var}(X_{t})\asymp 1/t$, which result in (20), after which the Lindeberg-Feller Theorem easily shows the central limit theorem in (21), because $X_{t}\leq n_{t}=O(1)$. If we impose the condition $n_{i}=n$, for $i\geq 1$, then we immediately obtain (18). For $\theta$ general, we reduce to the case $\theta=1$. This follows from noting that in (12), $\theta^{k_{t}}\asymp 1$ and $F_{t}(\theta)\asymp F_{t}(1)$ uniformly over $t\geq 1$, by (19), and so (20) still holds. For the almost sure convergence, modify the proof of Theorem 2.3.8 of [5]. ∎
The following theorem can get a Central Limit Theorem when $n_{t}$ is unbounded.
@@
Suppose $\theta=1$, $n_{t}=O(N_{t-1})$, and $\sum_{t=1}^{\infty}\frac{n_{t}}{N_{t-1}}=\infty$. Then, (21) holds.
@@
In order to prove limiting normality, we will verify Lyapanov’s Condition in [3] with $\delta=1$, by showing that
@@
${\mathbb{E}}([X_{t}-{\mathbb{E}}(X_{t})]^{3}\asymp{\rm Var}(X_{t})\asymp n_{t}/N_{t-1}.$
Hence,
${\rm Var}(K_{t})\asymp\sum_{s=1}^{t}{\mathbb{E}}([X_{s}-{\mathbb{E}}(X_{s})]^{3})\asymp\sum_{s=1}^{t}\frac{n_{s}}{N_{s-1}}$
and (22) holds by the second assumption. ∎

### v3

@@ 1 Introduction
A growth rule which received considerable attention in the literature is the following preferential attachment algorithm on permutations written as the product of cycles known as the Chinese restaurant process (CRP) [ABT, BMMU, Feng, GI, GW, GS, IKMW, CSP, Stark]. For $n\geq 1$, given a permutation of $n-1]:=\left\{1,2,\ldots,n-1\right\}$ which has been constructed at step n1n-1, element $n$ is either appended to the permutation as a new singleton cycle with probability $\frac{\theta}{\theta+n-1}$, or inserted into a random position within the existing permutation with probability equal to $\frac{n-1}{\theta+n-1}$. Let $S_{n}$ be the set of permutations on $[n]$. The permutation obtained at step $n$ has the Ewens distribution, which is the uniform distribution on $n!$ permutations in the case $\theta=1$, and, in general, has probability distribution $\theta^{|\pi|}/\theta_{(n)}$, $\pi\in S_{n}$, where $\theta_{(n)}:=\theta(\theta+1)\cdots(\theta+n-1)$ and $|\pi|$ is the number of cycles in $\pi$.
Let $\Pi_{n}$ be the random permutation of $[n]$ at the $n$th step of the CRP. The distribution of $\Pi_{n}$ is invariant under relabelling. For different orders the permutations are consistent, in the sense that $\Pi_{m}$, for $m<n$, can be derived from $\Pi_{n}$ by removing elements $m+1,\ldots,n$ from their cycles and deleting empty cycles if necessary.
@@
$\pi=(x_{1,1}\ldots x_{1,m_{1}}y_{1})(x_{2,1}\ldots x_{2,m_{2}}y_{2})\cdots(x_{\tau,1}\ldots x_{\tau,m_{\tau}}y_{\tau}),\quad\tau\geq 1,$
@@
which corresponds to the distribution of the sum of $n$ independent Bernoulli variables $K_{n}=\sum_{i=1}^{n}I_{i}$, $I_{i}\sim{\rm Bernoulli}(\theta/(i+\theta-1))$; see [ABT]. We have $I_{i}=1$ exactly when the element $i$ starts a new cycle. As $n\to\infty$,
$K_{n}\sim\theta\log n~~{\rm a.s.},~~~\frac{K_{n}-\theta\log n}{\sqrt{\theta\log n}}\stackrel{{\scriptstyle d}}{{\to}}{\rm N}(0,1),$
where $\stackrel{{\scriptstyle d}}{{\to}}$ denotes convergence in distribution.
In this paper, a CRP will be defined for multiset permutations and derived using the methods in [Stark]. As motivation, we note that multiset permutations are of fundamental interest in combinatorics [Stanley] and random multiset permutations have been studied previously [Feray].
@@ 2 Multiset permutations and their factorisations into
There are at least two different ways of defining cycles of multiset permutations [Knuth, Stanley]. Knuth’s way [Knuth] will be used in this paper. In this section, we summarise the discussion and results on multiset permutation factorisation taken from Section 5.1.2 of [Knuth] that is needed in this paper. In particular, we are going to use Theorem 1 below to factor multiset permutations. A multiset is like a set except that it can have repetitions of identical elements, for example
@@
The intercalation product $\pi_{1}\perp\pi_{2}$ of two multiset permutations $\pi_{1}$ and $\pi_{2}$ is defined in [CF, Foata] and is obtained by
@@
For example,
$\left(\begin{array}[]{c c c c c c}1&1&2&3&4\\ 3&1&4&1&2&\end{array}\right)\top\left(\begin{array}[]{c c c c c}1&2&4&4&4\\ 2&4&4&1&4\end{array}\right)$
@@
$=$ $\left(\begin{array}[]{c c c c c c c c c c}4&2&4&4&1&3&1&1&2&4\\ 2&4&4&1&3&1&1&2&4&4\end{array}\right)$ (7)
$=$ $\left(\begin{array}[]{c c c c c c c c c c}1&1&1&2&2&3&4&4&4&4\\ 3&1&2&4&4&1&2&4&1&4\end{array}\right)$ (10)
and so (1) is a cycle. For these general cycles, $(x_{1}x_{2}\ldots x_{n})$ is not always the same as $(x_{2}\ldots x_{n}x_{1})$ and it can happen that a cycle can be written as the intercalation product of other cycles, as we shall see at the end of this section.
Knuth [Knuth] proves the following factorisation of multiset permutations into cycles.
Theorem 1 ([Knuth]).
@@
$\pi=(x_{1,1}\ldots x_{1,m_{1}}y_{1})\top(x_{2,1}\ldots x_{2,m_{2}}y_{2})\top\cdots\top(x_{\tau,1}\ldots x_{\tau,n_{\tau}}y_{\tau}),\quad\tau\geq 1,$ (11)
where the following two conditions are satisfied:
@@
    $y_{i}<x_{i,j}{\rm\ for\ }1\leq j\leq m_{i},\,1\leq i\leq\tau$.
Note that (4) is not in the form (11). The factorisation of (1) corresponding to Theorem 1 is
$(3\ 1)\top(1)\top(2\ 4\ 2\ 4\ 4\ 1)\top(4)$
$=$ $\left(\begin{array}[]{c c}3&1\\ 1&3\end{array}\right)\top\left(\begin{array}[]{c}1\\ 1\end{array}\right)\top\left(\begin{array}[]{c c c c c c}2&4&2&4&4&1\\ 4&2&4&4&1&2\end{array}\right)\top\left(\begin{array}[]{c}4\\ 4\end{array}\right)$
$=$ $\left(\begin{array}[]{c c c c c c c c c c}3&1&1&2&4&2&4&4&1&4\\ 1&3&1&4&2&4&4&1&2&4\end{array}\right)$
$=$ $\left(\begin{array}[]{c c c c c c c c c c}1&1&1&2&2&3&4&4&4&4\\ 3&1&2&4&4&1&2&4&1&4\end{array}\right).$
@@ 3 The CRP for multiset permutations
where $N_{t}=n_{1}+\cdots+n_{t}$. Note that $S_{n}=S(\underbrace{{1,1,\ldots,1}}_{\text{$n$ times}}).$ The random multiset permutation $\tilde{\Pi}_{t}$ of our process at the $t$th step is an element of $S(n_{1},\ldots,n_{t})$, represented in accordance with (11). At step $t\geq 1$, there are $n_{t}$ copies of $t$ either to be inserted in the existing permutation $\tilde{\Pi}_{t-1}$ or to be put into singleton cycles. (It is not possible to put more than one copy of $t$ in a new cycle because of condition 2 of Theorem 1.) Suppose that $k_{t}\in\{0\}\cup[n_{t}]$ of the $n_{t}$ copies of $t$ start $k_{t}$ new singleton cycles containing $t$, and the $n_{t}-k_{t}$ other copies of $t$ are inserted into the existing permutation. The $n_{t}-k_{t}$ copies of $t$ can be inserted to the left of any of the $y_{i}$ or $x_{ij}$ in (11). A ‘stars and bars’ argument shows that the number of ways of inserting $n_{t}-k_{t}$ copies of $t$ to the left of an element in the multiset $S(n_{1},\ldots,n_{t-1})$ formatted like (11) is
@@
$F_{t}(1)=\sum_{k_{t}=0}^{n_{t}}\binom{N_{t-1}-1+n_{t}-k_{t}}{n_{t}-k_{t}}=\binom{N_{t}}{n_{t}}$. (16)
results from the well known identity
$\sum_{\rho=0}^{\nu}\binom{\mu+\rho}{\rho}=\binom{\mu+\nu+1}{\nu},\quad\nu,\mu\geq 0\ {\rm integers};$ (17)
see page 161 of [GKP]. Note that (16) is the total number of ways of extending the permutation from time t1t-1 to time $t$. At time $t\geq 1$, $k_{t}$ new singleton cycles containing $t$ are created and the other $n_{t}-k_{t}$ copies are inserted into the existing permutation in a specified way with probability $\theta^{k_{t}}/F_{t}(\theta)$. If $n_{t}=1$ for all $t$, then $F_{t}(\theta)=\theta+t-1$ and we recover the usual CRP. An induction argument on $t$ shows that all possible $\pi\in S(n_{1},\ldots,n_{t})$ can be constructed in this manner. This can also be seen from the identity
@@
We will use the method in [Stark]. An arborescence is a directed, rooted tree with edges oriented in agreement with paths from the root to the leaves. Our arborescence has vertex set $V=\bigcup_{s=0}^{t}\Pi(n_{1},\ldots,n_{s})$ and the root is $\emptyset$ (when $s=0$). The leaves of the arborescence are $\Pi(n_{1},\ldots,n_{t})$. Give each leaf $\pi\in\Pi(n_{1},\ldots,n_{t})$ the weight $\theta^{|\pi|}$, where $|\pi|$ is the number of cycles in $\pi$. There is a directed edge from $\pi_{1}\in\Pi(n_{1},\ldots,n_{s-1})$ to $\pi_{2}\in\Pi(n_{1},\ldots,n_{s})$, $s\in[t]$, if and only if, for some $k_{s}\in\{0\}\cup[n_{s}]$, $\pi_{2}$ is obtained from $\pi_{1}$ by inserting $n_{s}-k_{s}$ copies of $s$ to the left of elements of $\pi_{1}$, when $\pi_{1}$ is written as (1), and creating $k_{s}$ singleton cycles each containing $s$. For each $\pi\in V$, define $Q(\pi)$ to be the sum of the weights of all leaves $\pi^{\prime}\in\Pi(n_{1},\ldots,n_{t})$ for which there is a directed path from $\pi$ to $\pi^{\prime}$. For $\pi\in\Pi(n_{1},\ldots,n_{s})$, we have
@@
These are the transition probabilities of the CRP. Theorem 1 of [Stark] states that if we put transition probabilities (20) on the edges of the arborescence, and let the leaves be absorbing states, then we obtain a Markov chain whose probability of absorption on a given leaf is proportional to the weight of the leaf. By our choice of weight we are done. ∎
@@ 4 The distribution of the number of cycles
Let $X_{t}$ denote the number of new cycles added in moving from time t1t-1 to time $t$, $t\geq 1$, which is the choice of $k_{t}$ above. The $X_{t}$ are independent with distributions given by
@@
It is easily checked that (22) is (23) with $N=N_{t}$, $M=N_{t}-n_{t}=N_{t-1}$, $r=1$ and $\kappa=k+1$. The expectation of $Y$ is
@@
and the third centred moment (see [JKK, KK]) is
@@
It follows that the number of cycles $K_{t}=\sum_{s=1}^{t}X_{t}$ has expectation
${\mathbb{E}}(K_{t})=\sum_{s=1}^{t}\frac{n_{s}}{N_{s-1}+1}$
@@
Given sequences $a_{t}$ and $b_{t}$, we use the notation $a_{t}=O(b_{t})$ to mean $a_{t}\leq Cb_{t}$ for a constant $C>0$; $a_{t}\asymp b_{t}$ to mean $C_{1}b_{t}\leq a_{t}\leq C_{2}b_{t}$ for constants $C_{1}>0$, $C_{2}>0$; and $a_{t}\sim b_{t}$ to mean $\lim_{t\to\infty}a_{t}/b_{t}=1$. We also define $a_{t}=o(b_{t})$ to mean $\lim_{t\to\infty}a_{t}/b_{t}=0$. We will put the bound the $n_{t}=O(1)$ to give a central limit theorem for $K_{t}$ for all $\theta$.
@@
${\mathbb{E}}(K_{t})\asymp\log t,\quad{\rm Var}(K_{t})\asymp\log t,$ (29)
@@
$K_{n}\sim{\mathbb{E}}(K_{n})~~{\rm a.s.},~~~~~~\frac{K_{t}-{\mathbb{E}}(K_{t})}{\sqrt{{\rm Var}(K_{t})}}\stackrel{{\scriptstyle d}}{{\to}}{\rm N}(0,1).$ (30)
@@
First we prove the theorem for $\theta=1$. From (24), (25) and (28) we get ${\mathbb{E}}(X_{t})\asymp 1/t$ and ${\rm Var}(X_{t})\asymp 1/t$, which result in (29), after which the Lindeberg-Feller Theorem easily shows the central limit theorem in (30), because $X_{t}\leq n_{t}=O(1)$. If we impose the condition $n_{i}=n$, for $i\geq 1$, then we immediately obtain (27). For $\theta$ general, we reduce to the case $\theta=1$. This follows from noting that in (21), $\theta^{k_{t}}\asymp 1$ and $F_{t}(\theta)\asymp F_{t}(1)$ uniformly over $t\geq 1$, by (28), and so (29) still holds. For the almost sure convergence, modify the proof of (6.6) Theorem on page 44 of [durrett]. ∎
By the following theorem a Central Limit Theorem can be obtained even when $n_{t}$ is unbounded, which is not the case in Theorem 3.
@@
Suppose $\theta=1$, $n_{t}=O(N_{t-1})$. Then, (30) holds.
@@
In order to prove limiting normality, we will verify Lyapounov’s Condition in [Bill] with $\delta=1$, by showing that
@@
${\mathbb{E}}([X_{t}-{\mathbb{E}}(X_{t})]^{3}\asymp{\rm Var}(X_{t})\asymp n_{t}/N_{t-1},\quad t\geq 2.$
We have the lower bound
$\sum_{s=2}^{t}\frac{n_{s}}{N_{s-1}}$ $\geq$ $\sum_{s=2}^{t}\ln\left(1+\frac{n_{s}}{N_{s-1}}\right)$
$=$ $\sum_{s=2}^{t}\left(\ln(N_{s})-\ln(N_{s-1}\right)$
$=$ $\ln(N_{t})-\ln(N_{1})\to\infty,$
where the convergence to infinity follows from the assumption $n_{t}\geq 1$ for all $t\geq 1$. Hence,
${\rm Var}(K_{t})\asymp\sum_{s=1}^{t}{\mathbb{E}}([X_{s}-{\mathbb{E}}(X_{s})]^{3})\asymp\sum_{s=2}^{t}\frac{n_{s}}{N_{s-1}}\to\infty$
and (31) holds.
For the almost sure convergence, modify the proof of (6.6) Theorem on page 44 of [durrett] as in the proof of Theorem 3. ∎